Find the value of
[tan50 + tan25]/[1 – tan50.tan25]
All the angles are in degrees.

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

A = \frac{[tan50 + tan25]}{1-tan50.tan25}
tan(x+y) = \frac{[tanx + tany]}{1-tanx.tany}
A = \frac{[tan50 + tan25]}{1-tan50.tan25} = tan[50+25]
A = tan[45+30]
A = \frac{tan45 + tan30}{1-tan45.tan30}
A = \frac{1 + \frac{1}{\sqrt{3}}}{1-1.\frac{1}{\sqrt{3}}}
A = \frac{\sqrt{3} + 1}{\sqrt{3}-1}